OpenStudy (anonymous):

determine the slope m and y-intercept b of the given line: x=-1/2y + 3

OpenStudy (anonymous):

you can re-write this equations as: x-3=1/2 *y or y= 2x-6 1)slope - number in a front of x m=2 2) y- intercept means x=0 if x = 0, then y=-6 , (0,-6)

OpenStudy (anonymous):

I keep loosing signs... didn't see "-" in a front of 1/2.... easy to fix - change the sing : x-3=- 1/2 *y or y= -2x + 6 you can finish it now

OpenStudy (anonymous):

thanks inik you are awesome!!!